1. John kicks the ball, and the ball does a projectile motion with an angle of 53o to horizontal with an initial velocity of 10 m/s. What is (a) its maximum height, (b) its maximum distance, and (c) its hangtime?


  1. Find the range of an arrow that leaves the box at 50 m/s at ang angle of 50o above the horizontal?
1
Expert's answer
2021-10-11T08:54:53-0400

(a) The maximum height is


"h=\frac{(v\sin\theta)^2}{2g}=3.25\text{ m},"

the maximum distance is


"R=\frac{v^2\sin(2\theta)}{g}=9.8\text{ m}."

The hangtime is


"T=\frac{2v\sin\theta}g=1.63\text{ s}."

1. The range of the arrow is


"R=\frac{10^2\sin50°}{9.8}=7.82\text{ m}."